OpenStudy (anonymous):

Find the solution of this system of equations 2x-2y=-20 -2x+4y=14

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

4x+4y=-16 -8x-6y=-20

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

Lazyyyyy.. First one is (-3,13)

OpenStudy (anonymous):

Second is (-22,-26)